# gen2seg: Generative Models Enable Generalizable Instance Segmentation

### v1

@@ Abstract
By pretraining to synthesize coherent images from perturbed inputs, generative models inherently learn to understand object boundaries and scene compositions. How can we repurpose these generative representations for general-purpose perceptual organization? We finetune Stable Diffusion and MAE (encoder+decoder) for category-agnostic instance segmentation using our instance coloring loss exclusively on a narrow set of object types (indoor furnishings and cars). Surprisingly, our models exhibit strong zero-shot generalization, accurately segmenting objects of types and styles unseen in finetuning (and in many cases, MAE’s ImageNet-1K pretraining too). Our best-performing models closely approach the heavily supervised SAM when evaluated on unseen object types and styles, and outperform it when segmenting fine structures and ambiguous boundaries. In contrast, existing promptable segmentation architectures or discriminatively pretrained models fail to generalize. This suggests that generative models learn an inherent grouping mechanism that transfers across categories and domains, even without internet-scale pretraining. Code, pretrained models, and demos are available on our website.
@@ 1 Introduction
Generative models, particularly diffusion models [56, 23], have recently revolutionized image synthesis, achieving unprecedented levels of fidelity and diversity. This success has been driven by substantial industry investments and advances in model architecture, training strategies, and data availability. Concurrently, scene understanding tasks, such as segmentation or parsing object relationships, remain fundamental challenges in computer vision with widespread application across disciplines. Traditionally, generative modeling and scene understanding have been treated as distinct problems. However, inspired by [28, 40, 57], we ask if the powerful representations learned by generative models can also be leveraged to improve image understanding.
Large pretrained diffusion models, such as Stable Diffusion [52], have demonstrated the ability to capture intricate visual structures, enabling them to generate highly detailed and semantically coherent images. These models learn to map noisy inputs to clean, high-quality images through iterative refinement, implicitly acquiring an understanding of object boundaries, textures, and spatial relationships. We hypothesize that such models do not merely memorize pixel arrangements but also encode a latent structure that organizes pixels into distinct object instances. This implicit grouping is essential for generating realistic compositions but remains unobservable when evaluating generative performance alone.
We first explore finetuning pretrained diffusion models to segment object instances. Inspired by recent work [15], assuming a latent diffusion model with $N$ time steps $(0\dots N-1)$, we adapt it for single-step prediction by fixing $t=N-1$. We finetune the model end-to-end without noise. Since our loss is defined in pixel space, we decode the U-net’s output with the frozen VAE, and directly optimize the decoded image’s pixels. This approach yields a one-step, deterministic image segmenter that leverages Stable Diffusion’s large-scale generative pretraining. To tune our model, we design our instance coloring loss to encourage pixels belonging to the same object instance to share similar colors while encouraging distinct colors for different instances. Notably, we finetune using a mixture of two realistic synthetic datasets that contain only labels of indoor furnishings (i.e beds, tables, chairs) [51] and cars [4], but our model still generalizes to unseen object types and image styles. This suggests that the grouping mechanism learned during pretraining is not only robust but also transferable.
Crucially, our findings indicate that this capability arises not from Stable Diffusion’s internet-scale pretraining, but from the broader ability of generative models to understand pixel relationships while synthesizing realistic images. To validate this, we finetune a Masked Autoencoder [21] (MAE) encoder-decoder pair pretrained solely on the ImageNet-1K [12] dataset for masked token reconstruction with our segmentation loss and dataset. Like Stable Diffusion, it generalizes zero-shot to object types outside our finetuning dataset. Surprisingly, it also generalizes beyond the pretraining distribution to art, luggage x-rays, mythical creatures, and more.
$\overbrace{\makebox[221.93158pt]{}}^{\text{\small Ours}}$
Input DINO-B MAE-B MAE-H SD-Limited SD SAM
$\text{COCO}_{\text{exc}}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DRAM Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
EgoHOS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
iShape Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PIDRay Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
8
@@
Figure 2: To showcase the potential of generative models for instance segmentation, we highlight an example from each evaluation dataset where most or all of our models outperform SAM, despite never having seen masks of these object types. SAM often fails on fine structures (wires) or ambiguous boundaries (horses & carriage), leaving black regions where no object was detected. DINO-B also performs poorly, suggesting that generative pretraining (e.g., MAE, Stable Diffusion) learns strong priors for perceptual grouping. This is not meant to suggest that our method is inherently superior to SAM, which is trained on an annotated dataset several orders of magnitude larger. Instead, there are simply some objects which generative models are naturally better posed to segment.
Specifically, generalizing to "unseen" object types and styles means our models were never exposed to their segmentation masks. While our models may have seen images of them during pretraining (or via text conditioning, for Stable Diffusion), no segmentation supervision was provided. This highlights the remarkable ability of our models to generate high-quality masks for novel object types. MAE, in particular, was pretrained only on unlabeled ImageNet-1K, which almost certainly lacks any images from domains like impressionist art or luggage X-rays, demonstrating true zero-shot generalization as it never saw similar pixels or masks in any stage of training. This suggests learning to generate image pixels inherently leads to acquiring object instance groupings, reinforcing the idea that generative models can serve dual purposes beyond image synthesis.
Our results open new avenues for leveraging the massive investment in generative AI for tasks that go beyond visual content creation. By demonstrating that generative models inherently learn transferable representations for object segmentation, we provide a compelling case for their broader application in real-world perception. We hope utilizing the generative representations learned from image synthesis can pave the way for more efficient, generalizable, and data-efficient segmentation methods, enabling advances in fields where detailed scene understanding is critical, such as robotics, medical imaging, and autonomous systems.
@@ 2.1 Generative Models for Perception
Generative models, originally developed for image synthesis, have increasingly been adapted for perception tasks in computer vision. A longstanding viewpoint in the field (dating back to Hinton’s early work) [22] posits that learning to generate data can aid in recognizing it. Early work on GANs [17] evaluated whether representations learned by generating images [48] or videos [60] transferred well to image classification or action recognition, respectively, but performance was always far below discriminatively pretrained models. Concurrent works also utilized inpainting [45] and colorization [31, 65, 66] as pretext tasks for representation learning, but these were subsequently surpassed by discriminative pretraining [43, 16, 20].
Recent advancements have demonstrated the efficacy of diffusion models [56, 23] in various visual tasks. A key advantage of recent diffusion models, such as Stable Diffusion, is the sheer scale of their pretraining; learning from over 2 billion images [52, 53] has the potential to outscale existing discriminatively trained models. Their large-scale generative pretraining has since been transferred to many perceptual tasks, such as 3D reconstruction [40, 47, 61], semantic [2, 35, 27, 58] and amodal segmentation [44, 9, 50], monocular depth [28, 50, 67], surface normals [14], optical flow [50], correspondence [57], and classification [32]. Recent work [15] also showed how to adapt these multistep models for single-step prediction, enabling task-specific losses in pixel space. Other works have shown that depth, normals, albedo, and segmentation can emerge (albeit with low quality) from generative models [3, 42, 26] without finetuning, suggesting similar representations may emerge from the data alone [13, 24]. However, applying generative models to instance segmentation has remained relatively unexplored.
@@
Figure 3: Our models assign similar colors to compositionally related parts of a scene. Vader’s mask and body (top), or the bowties and shirts (bottom) are separated by subtly different hues, while distinct colors partition unrelated parts such as his leg and the poles (top), or the dogs and text (bottom). This emerges without any part-level supervision, suggesting generative models learn hierarchical scene representations. More samples are provided in the high resolution figures in the appendix.
A parallel line of work focused on representation learning, Masked Autoencoders (MAE) [21], has achieved SOTA in a large number of visual tasks by pretraining to reconstruct masked image tokens, and then fine-tuning for discriminative tasks. However, common practice is to discard the decoder, which contains its pixel-perfect generative features, prior to finetuning. Prior works have shown the decoder’s generalization in solving visual prompts [1]. However, [1] focuses on MAE’s ability to generate masks via prompting, while we finetune, and find it generalizes to objects whose masks are unseen in finetuning.
@@ 2.2 Instance Segmentation
Instance segmentation has evolved from traditional two-stage methods [19, 5, 8], which first generate object proposals and then identify pixel groupings, to end-to-end approaches that directly predict masks using learned query embeddings [6, 33, 10]. While these direct methods eliminate the proposal stage, they still require extensive, class-specific annotated data, limiting their generalization.
Recent advancements have introduced category-agnostic approaches, enabling models to segment objects without prior knowledge of their classes. The most well known of these is SAM [29], which learns zero-shot promptable segmentation by finetuning an MAE backbone and mask decoder on the massive SA-1B dataset. SAM and its recent successor SAM2 [49] represent a breakthrough in obtaining general, category-agnostic masks without per-dataset training. SAM’s zero-shot generalization provides a point of comparison for our work. However, unlike SAM, which was trained on a massive supervised dataset, our approach aims to leverage generative knowledge to achieve broad instance segmentation, potentially reducing the need for extensive manual mask annotations. Furthermore, our strongest model is trained for only 29 hours on 4 RTX6000 Ada (48GB) GPUs on less than 87,000 images and 3.7 million masks of only select categories, while SAM was trained for 68 hours on 256 A100 (80GB) GPUs on 11 million images and 1.1 billion masks of all types.
@@ 3 Method
We aim to adapt pretrained generative models, such as Stable Diffusion or MAE, to perform category agnostic instance segmentation. Instance segmentation, in particular, is a pixel level task, at which generative models are naturally primed to excel. Specifically, we hypothesize they inherently learn to understand object boundaries and groupings because they must synthesize the objects’ core structure and boundaries themselves. In contrast, most SOTA models, such as SAM [29], extract features using an encoder that discards low-level details. As a result, they must learn mask predictors or feature pyramid networks from scratch to gradually upsample these features to higher resolutions [37, 34].
@@ 3.1 Instance Segmentation as Image to Image Translation
Recent instance segmentation models predict sets of binary masks, each representing an object instance [6, 10]. However, it is not obvious how to enable generative models, designed to map from $\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times 3}$, to easily decode to this style. Drawing inspiration from work on image-to-image translation [25, 68], we encode our ground truth as an RGB image with a unique color for each instance and black color for the background. We find that both the Stable Diffusion VAE and MAE models can decode these masks with effectively no loss in quality. Thus, we finetune our model by forwarding our RGB image (or its latent) without adding noise or masking, decoding the output, and optimizing in pixel space with respect to the RGB ground truth.
Unfortunately, one cannot simply assign each ground truth mask to a color and train the model to regress it since there are many possible accurate color combinations. Thus, we propose our instance coloring loss based on two key properties of an RGB segmentation mask. First, all the pixels within a mask should have no variance. Second, the color of a mask should not be repeated anywhere outside the mask. By emphasizing these two properties, we can learn a model of instance segmentation without needing to enforce specific colors for object masks. Simply, our finetuned model should ensure each instance is assigned a unique color that is consistent across its pixels.
@@
where $p_{j,c}$ is the $c$’th channel of the predicted color at pixel $j$. We force the background mask’s mean to be black ($\mu_{0,c}=0$) to follow standard convention and distinguish it from objects.
@@
$\mathcal{L}_{\mathrm{var}}=\sum_{i=0}^{n}\frac{1}{|S_{i}|}\sum_{j\in S_{i}}\sum_{c=0}^{2}L_{s}(p_{j,c},\mu_{i,c})\$ (3)
@@
3.2 Promptable Instance Segmentation
So far, our finetuned model assigns a color $F(x,y)$ to each pixel. Since pixels belonging to the same object instance are encouraged to have similar colors, we develop a simple method to “prompt” the feature map $F$ for binary masks (similar to SAM [29]). For each prompt point $p\in\mathbb{R}^{2}$ which contains the $x$ and $y$ location of the prompt pixel, we first compute a per-pixel weight $w(x,y)$ as a Gaussian function with mean $p$ and standard deviation $0.01\times(W,H)$, where $W$ and $H$ are the width and height of $F$. We use $w$ to calculate the query vector $q_{p}\in\mathbb{R}^{3}$ as a weighted average of $F$ with high weights for pixels near $p$ as $q_{p}=\frac{\sum_{x,y}w(x,y)\,F(x,y)}{\sum_{x,y}w(x,y)}$.

### v2

@@ Abstract
By pretraining to synthesize coherent images from perturbed inputs, generative models inherently learn to understand object boundaries and scene compositions. How can we repurpose these generative representations for general-purpose perceptual organization? We finetune Stable Diffusion and MAE (encoder+decoder) for category-agnostic instance segmentation using our instance coloring loss exclusively on a narrow set of object types (indoor furnishings and cars). Surprisingly, our models exhibit strong zero-shot generalization, accurately segmenting objects of types and styles unseen in finetuning. This holds even for MAE, which is pretrained on unlabeled ImageNet-1K only. When evaluated on unseen object types and styles, our best-performing models closely approach the heavily supervised SAM, and outperform it when segmenting fine structures and ambiguous boundaries. In contrast, existing promptable segmentation architectures or discriminatively pretrained models fail to generalize. This suggests that generative models learn an inherent grouping mechanism that transfers across categories and domains, even without internet-scale pretraining. Code, pretrained models, and demos are available on our website.
@@ 1 Introduction
Humans learn to carve the visual world into discrete, persistent objects from limited experience. A toddler who has mostly handled cups, chairs, and toys at home can still recognize the zebra, giraffe, and lion on their first trip to the zoo as distinct objects, even without knowing what they are called. This ability extends to abstract depictions: people can understand line drawings and artworks even without prior exposure to them (Hochberg & Brooks, 1962). These observations suggest human vision acquires general, transferable mechanisms for grouping pixels into objects.
Modern vision systems often show ”zero-shot” transfer to new datasets without domain-specific fine-tuning, but typically rely on broad labeled datasets that attempt to cover many object categories and styles. We ask a different question through the lens of instance segmentation: can a model learn from only a very narrow slice of the visual world and still generalize to unseen object types and image styles? We focus on a stricter “zero-shot” setting, where we explore how the model performs on object types it has never seen a mask for.
We hypothesize that generative models are particularly well posed to succeed at this task. Because they learn to synthesize scenes from minimal cues (e.g., a text prompt or a corrupted image), they must learn to implicitly represent the parts that make up the image. This allows generative models to produce scene compositions unlike anything seen in pretraining data, such as “a Van-Gogh style painting of a panda driving a car.”
We introduce a simple finetuning method that taps into these generative priors. We first experiment with a Masked Autoencoder (MAE; encoder+decoder) pretrained on ImageNet-1K only end-to-end using only mask supervision from two narrow synthetic domains (indoor furnishings and cars) to generate object instance groupings. Despite this limited supervision, our model is able to generalize to new object types unseen in finetuning, such as people and animals, or novel image styles, such as impressionist art and x-rays. When we experiment with limiting the diversity or complexity of our finetuning dataset, this generalization persists, suggesting it is due to the generative prior.
To explore the effects of internet-scale pretraining, we apply our finetuning method to Stable Diffusion 2. Without seeing any labeled masks from the evaluation categories, our finetuned Stable Diffusion achieves performance comparable to SAM across five datasets of distinct domains. Beyond generalization, our models consistently produce crisper boundaries than SAM (e.g., on BSDS500), a behavior that persists even when finetuned on datasets with polygonal edges (e.g., COCO). They also excel at segmenting fine structures (e.g., on iShape) and exhibit object-part compositionality, despite never receiving part-level supervision. We hypothesize these behaviors arise directly from generative pretraining, which must model fine edges, delicate structures, and part–whole relationships to synthesize detailed scenes.
Together, our findings argue that generative pretraining encodes an inherent grouping mechanism that extends beyond both the object types and the image styles seen during finetuning. We hope utilizing the generative representations learned from image synthesis can pave the way for more generalizable and human-like perception, enabling advances in fields where detailed scene understanding is critical, such as robotics, medical imaging, and autonomous systems.
$\overbrace{\makebox[166.94296pt]{}}^{\text{Ours}}$
Input DINO-B MAE-B MAE-H SD (10 Classes) SD SAM
$\text{COCO}_{\text{exc}}$
DRAM
EgoHOS
iShape
PIDRay
@@
Figure 2: To showcase the potential of generative models for instance segmentation, we highlight an example from each evaluation dataset where most or all of our models outperform SAM, despite never having seen masks of these object types. SAM often fails on fine structures (wires) or ambiguous boundaries (horses & carriage), leaving black regions where no object was detected. DINO-B also performs poorly, suggesting that generative pretraining (e.g., MAE, Stable Diffusion) learns strong priors for perceptual grouping.
@@ 2.1 Generative Models for Perception
Generative models, originally developed for image synthesis, have increasingly been adapted for perception tasks in computer vision. A longstanding viewpoint in the field (dating back to Hinton’s early work) (Hinton, 2007) posits that learning to generate data can aid in recognizing it. Early work on GANs (Goodfellow et al., 2014) evaluated whether representations learned by generating images (Radford et al., 2015) or videos (Vondrick et al., 2016) transferred well to image classification or action recognition, respectively, but performance was always far below discriminatively pretrained models. Some works utilized inpainting (Pathak et al., 2016) and colorization (Larsson et al., 2017; Zhang et al., 2016; 2017) as pretext tasks for representation learning, but these were subsequently surpassed by discriminative pretraining (Noroozi et al., 2017; Gidaris et al., 2018; He et al., 2020).
Recent advancements have demonstrated the efficacy of diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) in various visual tasks. A key advantage of recent diffusion models is the sheer scale of their pretraining; learning from over 2 billion images (Schuhmann et al., 2022) has the potential to outscale existing discriminatively trained models. Their large-scale generative pretraining has since been transferred to many perceptual tasks, such as 3D reconstruction (Liu et al., 2023b; Poole et al., 2022; Wang et al., 2023a), semantic (Baranchuk et al., 2021; Li et al., 2023b; Kawano & Aoki, 2024; Tian et al., 2024) and amodal segmentation (Ozguroglu et al., 2024; Chen et al., 2024), monocular depth (Ke et al., 2024; Zhao et al., 2023), surface normals (Fu et al., 2024), optical flow (Ravishankar et al., 2024), correspondence (Tang et al., 2023), and classification (Li et al., 2023a). Other works have shown that depth, normals, albedo, and segmentation can emerge (albeit with low quality) from generative models (Bhattad et al., 2023; Namekata et al., 2024; Karmann & Urfalioglu, 2024) without finetuning, suggesting similar representations may emerge from the data alone (Dravid et al., 2023; Huh et al., 2024). While there are some prior works that finetune diffusion models for instance segmentation (Fan et al., 2024; Zhao et al., 2025), these works focus on building competitive instance segmenters using large scale data, while we explore generalization through the lens of instance segmentation.
@@
Figure 3: Our models assign similar colors to compositionally related parts of a scene. Vader’s mask and body (top), or the bowties and shirts (bottom) are separated by subtly different hues, while distinct colors partition unrelated parts such as his leg and the poles (top), or the dogs and text (bottom). This emerges without any part-level supervision, suggesting generative models learn hierarchical scene representations. More samples are provided in Figures 13 to 19.
A parallel line of work explores representation learning through Masked Autoencoders (MAE) (He et al., 2022), which achieve state-of-the-art performance across many visual tasks by pretraining to reconstruct masked image tokens before fine-tuning on discriminative objectives. A common practice, however, is to discard the decoder despite it containing rich pixel-level generative features prior to fine-tuning. Recent work (Bar et al., 2022) has demonstrated the encoder and decoder’s joint capacity to generalize through visual prompting, generating masks without explicit supervision. In contrast, our approach end-to-end fine-tunes the encoder and decoder and shows that it can generalize to objects whose masks were not observed during fine-tuning.
@@ 2.2 Instance Segmentation
Recent advancements have introduced category-agnostic instance segmentation, enabling models to segment objects without prior knowledge of their classes. The most well known of these is SAM (Kirillov et al., 2023), which learns zero-shot promptable segmentation by finetuning an MAE backbone and mask decoder on the massive SA-1B dataset. SAM and its recent successor SAM2 (Ravi et al., 2024) represent a breakthrough in obtaining general, category-agnostic masks without per-dataset training. However, unlike SAM, which was trained on a massive labeled dataset, our approach aims to leverage generative knowledge to achieve broad instance segmentation. Furthermore, our strongest model is trained for only 29 hours on four RTX6000 Ada (48GB) GPUs on less than 87,000 images and 3.7 million masks of only select categories, while SAM was trained for 68 hours on 256 A100 (80GB) GPUs on 11 million images and 1.1 billion masks of many aspects of the visual world.
Recent works (Wang et al., 2024; 2023b; 2022) train instance segmenters from pseudo-masks automatically derived from contrastive self-supervised features via normalized cuts (Shi & Malik, 2000), or similarity and thresholding. While both our work and theirs use less labelled data compared to prior works (Kirillov et al., 2023), our goals differ. They leverage vast, diverse unlabeled images to synthesize pseudo-labels for downstream finetuning, to reduce reliance on manually annotated labels. In contrast, we use manually annotated labels, but deliberately restrict annotation diversity, to build models that generalize to many novel object types when finetuned on very few. Additionally, the noisy pseudo-labels require iterative retraining, which takes 2–15x longer than our single-pass fine-tuning. Many of the findings from our work, such as our zero-shot generalization and our robustness to noisy labels, provide a path to eventually combine these two directions in future work.
@@ 3 Method
We aim to adapt pretrained generative models, such as Stable Diffusion or MAE, to perform category agnostic instance segmentation. Instance segmentation, in particular, is a pixel level task, at which generative models are naturally primed to excel. Specifically, we hypothesize they inherently learn to understand object boundaries and groupings because they must synthesize the objects’ core structure and boundaries themselves. In contrast, most SOTA models, such as SAM (Kirillov et al., 2023), extract features using an encoder that discards low-level details. As a result, they must learn mask predictors or feature pyramid networks from scratch to gradually upsample these features to higher resolutions (Lin et al., 2017; Li et al., 2022).
@@ 3.1 Instance Segmentation as Image to Image Translation
Recent instance segmentation models predict sets of binary masks, each representing an object instance (Carion et al., 2020; Cheng et al., 2022). However, it is not obvious how to enable generative models, designed to map from $\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times 3}$, to easily decode to this style. Drawing inspiration from work on image-to-image translation (Isola et al., 2017; Zhu et al., 2017), we encode our ground truth as an RGB image with a unique color for each instance and black color for the background. We find that both the Stable Diffusion VAE and MAE models can decode these masks with effectively no loss in quality. Thus, we finetune our model by forwarding our RGB image (or its latent) without adding noise or masking, decoding the output, and optimizing in pixel space with respect to the RGB ground truth.
Unfortunately, one cannot simply assign each ground truth mask to a color and train the model to regress it since there are many possible accurate color assignments. Thus, we propose our instance coloring loss based on two key properties of an RGB segmentation mask. First, the color of all pixels within a mask should have low variance. Second, the color of a mask should not be predicted anywhere outside the mask. By emphasizing these two properties, we can learn a model of instance segmentation without needing to enforce specific colors for object masks. Simply, our finetuned model should ensure each instance is assigned a unique color that is consistent across its pixels. Our finetuning strategy is partially inspired by prior works (De Brabandere et al., 2017; Kong & Fowlkes, 2018) and thus we claim no novelty on it. We designed it to be simple, intuitive, and agnostic to model architecture with fast inference time.
@@
where $p_{j,c}$ is channel $c$ of the predicted color at pixel $j$. We force the background mask’s mean to be black ($\mu_{0,c}=0$) to follow standard convention and distinguish it from objects.
@@
$\mathcal{L}_{\mathrm{var}}=\sum_{i=0}^{n}\frac{1}{|S_{i}|}\sum_{j\in S_{i}}\sum_{c=0}^{2}L_{s}(p_{j,c},\mu_{i,c})\\$ (3)
@@
3.2 Point-Promptable Instance Segmentation
So far, our finetuned model assigns a color $F(x,y)=p_{j}$ to each pixel where $(x,y)$ is the location of pixel $j$. Since pixels belonging to the same object instance are encouraged to have similar colors, we develop a simple method to “point-prompt” the feature map $F$ for binary masks (similar to SAM (Kirillov et al., 2023)). We intentionally opt not to train a separate mask decoder to showcase that our model’s output features truly represent object instance shapes. Additionally, existing mask decoders use specialized architectures designed to upsample from low resolutions to the original image size, while our features are at the same resolution as the original image. Our simple prompting method can be viewed as analogous to similar evaluation methods in representation learning used to show that features truly represent the desired task, such as nearest-neighbor classification, semantic segmentation, or tracking (Caron et al., 2021). One can potentially improve our results by training a promptable high-resolution mask decoder on top of our features, which we leave for future work.
For each prompt point $p\in\mathbb{R}^{2}$ which contains the $x$ and $y$ location of the prompt pixel, we calculate the query vector $q_{p}\in\mathbb{R}^{3}$ as a Gaussian weighted average of the predicted color at the neighborhood of $p$ by $q_{p}=\frac{\sum_{x,y}w(x,y)\,F(x,y)}{\sum_{x,y}w(x,y)}$ where $w(x,y)$ is a Gaussian function with mean $p$ and standard deviation $0.01(W,H)$ and $W$ and $H$ are the width and height of $F$.